# Concept Drift Guided LayerNorm Tuning for Efficient Multimodal Metaphor Identification

### v1

@@ 1. Introduction
The multimodal metaphorical meme dataset MET-Meme (Xu et al., 2022) has played a crucial role in advancing this field by catalyzing methodological innovations. The methods (He et al., 2024b, a; Wang et al., 2024) mainly focus on fine-grained feature alignment and fusion, resulting in suboptimal performance, as they often fail to fully leverage the characteristics of multimodal metaphors, overlooking the implied meanings and unconventional expressions that frequently appear in figurative tasks. On the other hand, methods (Xu et al., 2024a; Zhang et al., 2024; Xu et al., 2024b) that focus on generative knowledge expansion help bridge the gap between literal and figurative tasks by using the generative information from large language models (LLMs) or text-to-image models. Although using LLMs or text-to-image models has demonstrated significant advantages in multimodal metaphor identification tasks, they still face issues with high computational overhead and large GPU memory usage during training, even when these methods employ some general Parameter-Efficient Fine-Tuning (PEFT) techniques like LoRA (Hu et al., 2021). Recently, the novel methods (Zhou et al., 2023; Yu et al., 2024) that only fine-tune the LayerNorm layers of language models for cross-modal feature extraction and fusion, have demonstrated outstanding efficiency: achieving good performance by fine-tuning less than 4% of the total parameters. However, LayerNorm tuning of pretrained language models for feature extraction and fusion has remained unexplored in multimodal metaphor identification tasks due to their suboptimal performance when handling non-sequential data, such as images.
@@ 2.1. Multimodal Metaphor Understanding
The study of metaphors has expanded into the multimodal domain, exploring how this cognitive phenomenon manifests in various modalities, particularly in vision. Several datasets have been created to facilitate research in this area. MultiMET (Zhang et al., 2021) is a multimodal metaphor dataset containing text-image pairs with annotations for metaphor occurrence, domain relations, and sentiments. MetaCLUE (Akula et al., 2023) is another dataset including four understanding tasks for multimodal metaphor. The MET-Meme (Xu et al., 2022) dataset provides a valuable resource for studying multimodal metaphors in the context of memes. MultiCMET (Zhang et al., 2023) considers metaphors from different cultures, extending the work of MultiMET to Chinese. IRFL (Yosef et al., 2023) is a relevant dataset including metaphor and idiom data, which focuses on the multiple choices task. MemeCap (Hwang and Shwartz, 2023) includes annotations such as image captions, titles, and metaphorical captions to explore the automatic interpretation and understanding of multimodal metaphors. (R et al., 2024) introduces a task of video metaphor description, addressing the gap in Vision-Language (VL) models’ ability to understand metaphors in video. Various models have been proposed for detecting and understanding multimodal metaphors. Early work focused on visual atypicality, which can be seen as a special kind of multimodal metaphor, such as (Guo et al., 2021) models contextual compatibility to detect persuasive atypicality. The subsequent works primarily use MET-Meme as the benchmark, with early studies (He et al., 2024b, a; Wang et al., 2024) focusing on fine-grained multimodal feature fusion or alignment. Recently, there is a growing trend to leverage pre-trained generative models to expand the knowledge for metaphor understanding. For example, C4MMD (Xu et al., 2024a) is a compact framework that uses a Chain-of-Thought (CoT) method to extract and integrate knowledge from Multimodal Large Language Models (MLLMs) into smaller models for multimodal metaphor detection. CAMEL (Zhang et al., 2024) leverages the captions generated by multimodal language models as a bridge to capture the implicitly established metaphor alignment. And MMMC (Xu et al., 2024b) utilizes a text-conditioned generative adversarial network to generate visual characteristics based on the linguistic attributes of metaphorical concepts, thereby capturing more comprehensive visual associations.
@@ 3.2. Concept Drift and Feature Fusion
After obtaining $E^{S}$, we fuse the three embeddings—$E^{I}$, $E^{S}$, and $E^{T}$—into a unified feature vector through concatenation and feedforward neural network (FFN) operations. The vector obtained after the concatenate operation is denoted as $E^{Meme}\in\mathbb{R}^{N\times 3d_{c}}$, and the vector obtained after the FFN is denoted as $F\in\mathbb{R}^{N\times d_{g}}$. This process is formalized as follows:
@@ 4. Experiments
MultiCMET (Zhang et al., 2023) - - - 78.79 - - - 85.66
VIEMF (He et al., 2024b) - - - 83.92 - - - 84.87
SC-Net (He et al., 2024a) - - - 86.85 - - - 87.50
C4MMD (Xu et al., 2024a) - - - 82.44 - - - 87.70
MET-Meme (Xu et al., 2022) 27.68 67.25 38.56 82.39 29.82 74.48 39.84 83.33
$\text{M}^{\text{3}}\text{F}\_\text{add}$ (Wang et al., 2024) 31.82 72.66 43.07 85.11 30.47 76.17 44.40 83.98
$\text{M}^{\text{3}}\text{F}\_\text{cat}$ (Wang et al., 2024) 32.36 72.72 44.04 85.89 29.82 74.09 44.10 83.20
@@ 4.2. Evaluation Metrics
[…]
@@ 4.4. Main Results
In this section, we compared our method with existing multimodal metaphor detection works including MET-Meme (Xu et al., 2022), MultiCMET (Zhang et al., 2023), VIEMF (He et al., 2024b), SC-Net (He et al., 2024a), C4MMD (Xu et al., 2024a), $\text{M}^{\text{3}}\text{F}$ (Wang et al., 2024) and CAMEL (Zhang et al., 2024) on Sentiment Analysis (SA), Offensiveness Detection (OD), Intention Detection (ID), and Metaphor Identification (MI), four tasks of MET-Meme. Among these, Multi-CMET, VIEMF, SC-Net and C4MMD solely focus on the MI task. The methods $\text{M}^{\text{3}}\text{F}$ and CAMEL are not only proposed for MI task but also for SA, ID and OD tasks in MET-Meme. At the same time, both $\text{M}^{\text{3}}\text{F}$ and CAMEL have proposed two model variants, and we have included these variants in our comparison as well. As shown in Table 1, we report the performance results of the methods to be compared and our two model settings. One of our settings is without SLERP called $\text{CDGLT}_{Vanilla}$ and the other with SLERP just called CDGLT. Although our approach is primarily focused on the multimodal metaphor identification task, we found that our architecture, with specifically training on other tasks, also achieved performance superior to or on par with the state-of-the-art. It is noted that CDGLT achieves the highest performance of accuracy and Weighted F1 score in MI and SA tasks. On the other hand, $\text{CDGLT}_{Vanilla}$ demonstrates the complementary situations, achieving the best performance on the other two tasks: ID and OD. This suggests that incorporating Concept Drift embedding enhances the model’s ability to capture non-literal understanding tasks, such as MI. In contrast, tasks like ID and OD appear to benefit more from the vanilla model without SLERP, suggesting that these tasks demand more straightforward information and less divergence. This implies that the benefits of SLERP are not universally applicable and are instead highly dependent on the specific task. Our ablation analysis in section 4.5.1 will further clarify this situation. In addition, it can be observed that even with the use of Concept Drift, the CDGLT model achieves excellent performance in nearly all tasks, outperforming previous methods except in the accuracy of the OD task. Our framework does not utilize the generative information of LLM, surpassing methods like C4MMD and CAMEL, which use pre-trained language models to generate explanations of metaphors or implicit information as supplementary knowledge.
@@ 4.5. Further Analysis
In this section, we conducted a series of ablation experiments on the composition of $E^{Meme}$ and observed some interesting phenomena. As shown in Table 2, the results compare the input of the FFN adapter, $E^{Meme}$, with different settings. First, comparing “$E^{I}$ only” versus “$E^{T}$ only”, the results of the two experiments show only minor differences in the OD and ID tasks, while there are significant differences in the SA and MI tasks. Combined with other experiments, it is clear that the image modality is indispensable for improving performance across almost all tasks. This aligns with our intuition, as memes themselves use images as carriers of all information, including text. However, interestingly, for ID and OD tasks, just the OCR text from memes, although not the highest, still performs quite well. At the same time, this also suggests that, compared to the other two tasks, there may be more compatible commonalities between SA and MI, and the same applies to OD and ID.
The two experiments, “$E^{S}$ only” and “concat($E^{I}$, $E^{T}$)”, reveal that on the SA and MI tasks, the performance of “concat($E^{I}$, $E^{T}$)” closely aligns with that of “$E^{I}$ only”. Meanwhile, the performance of “$E^{S}$ only” just falls between that of “$E^{I}$ only” and “$E^{T}$ only”. This may suggest that the concatenated feature vector aggregates information from both embeddings. However, the $E^{S}$ obtained through the SLERP operation seems to represent an intermediate state transitioning from $E^{I}$ to drift towards $E^{T}$.
@@
[…]
However, for the ID and OD tasks, there is a phenomenon where performing concatenation between any two embeddings of $E^{I}$, $E^{T}$, and $E^{S}$, results in improved performance on the ID or OD tasks compared to using $E^{I}$ or $E^{T}$ alone. Specifically, “concat($E^{I}$, $E^{S}$)” improves performance on the OD task compared to using “$E^{I}$ only”, while “concat($E^{T}$, $E^{S}$)” enhances performance on the ID task compared to “$E^{T}$ only”. Finally, “concat($E^{I}$, $E^{T}$)” leads to improvements in both the ID and OD tasks. Similar to “concat($E^{I}$, $E^{T}$)”, using “$E^{S}$ only” also leads to improvements in both the ID and OD tasks compared to using “$E^{I}$ only” and “$E^{T}$ only”. Additionally, from the results of the group “concat($E^{I}$, $E^{S}$, $E^{T}$)” and “concat($E^{I}$, $E^{T}$)”, it can be observed that while adding $E^{S}$ improves the performance of SA and MI tasks, it also leads to a decline in the performance of ID and OD tasks. Furthermore, in the ID task, “concat($E^{I}$, $E^{S}$, $E^{T}$)” performs worse than “concat($E^{T}$, $E^{S}$)”, and in the OD task, it performs worse than “concat($E^{I}$, $E^{S}$)”. Therefore, concatenating any two embeddings with a third embedding will lead to a decrease in performance on both the OD and ID tasks. This could mean that the diverging information might be redundant for both OD and ID tasks, or even become noise that interferes with performance.
Comparison of the two experiments “concat($E^{I}$, $E^{S}$)” and “concat($E^{I}$, $E^{T}$)”, it shows that although replacing $E^{T}$ with $E^{S}$ leads to poorer performance on other tasks, it results in an improvement on the MI task. Additionally, by comparing “concat($E^{I}$, $E^{S}$)” with “$E^{I}$ only” or “concat($E^{I}$, $E^{S}$, $E^{T}$)” with ’concat($E^{I}$, $E^{T}$)’, it is found that adding $E^{S}$ also contributes positively to the performance on the MI task. At the same time, this may suggest that divergent guidance, such as Concept Drift, indeed contributes to figurative tasks.
@@
Table 3. Prompt Ablations on Metaphor Identification (MI) task. [$E^{T}$, $E^{I}$] and [$E^{T}$, $E^{S}$, $E^{I}$] represent sequences of embeddings. F is the feature from concatenating $E^{T}$, $E^{S}$, and $E^{I}$ into an embedding three times the hidden dimension, then processed by a FFN adapter. The “length” column represents the number of embeddings of the sequence. Instruction A: “Providing you with the memes, now identify the metaphorical ones:”; Instruction B: “Metaphor Detection (literal — metaphor):”; Instruction C: “Metaphor Detection:”. Trainable and Frozen Vectors refer to prompts constructed using trainable and frozen Xavier initialization vectors, respectively.
@@
In this section, we conduct a series of ablation studies on the prompt construction method we proposed for LN Tuning. Experimental results on the MI task is shown in Table 3 demonstrating that, compared to other strategies, our method has significant advantages on the multimodal metaphor identification task based on meme data. The method of LN Tuning with a pretrained language model is commonly used for feature extraction or feature fusion of sequence information. Therefore, the most intuitive and common approach is to directly construct an embedding sequence as the input to the language model, such as: $[E^{T},E^{I}]\in\mathbb{R}^{N\times 2\times d_{g}}$ and $[E^{T},E^{S},E^{I}]\in\mathbb{R}^{N\times 3\times d_{g}}$. However, we found that by first fusing the features using an FFN adapter and then directly using the fused features $F\in\mathbb{R}^{N\times 1\times d_{g}}$ as input, we achieved better performance, even though the input sequence length was 1 at that point. Since the attention mechanism exhibits stronger capabilities when processing sequences, we additionally construct an embedding sequence as a prompt based on $F$. We tried three strategies: the first one uses the word embedding layer of GPT2 to transform fixed instruction words into an embedding sequence; the second one uses Xavier initialization to create some learnable vectors; the third one freezes the Xavier-initialized vectors. In each method, $F$ is placed at the end of the sequence. For the first strategy, we tried three instructions: “Providing you with the memes, now identify the metaphorical ones:”, “Metaphor Detection (literal — metaphor):” and “Metaphor Detection:”. After tokenization, they consist of 14 tokens, 10 tokens and 5 tokens, respectively. For both “Trainable vectors” and “Frozen vectors”, we also used sequence lengths of 14, 10 and 5, respectively. The experimental results show that “frozen vectors” perform significantly better than “trainable vectors”. Comparing the experimental results of the Frozen Xavier initialization vectors method under different length settings may indicate that using prompts that are too long or too short could potentially lead to a decline in performance. However, we intuitively believe that when using word instructions as prompts, the performance will be significantly affected by the semantic content, and this impact is difficult to capture. At the same time, we noticed that in our experiments, the method of frozen vectors and word instructions performed similarly. To ensure the generality of the method, our framework ultimately chose to construct prompts by adopting frozen vectors.

### v3

@@ 1. Introduction
The multimodal metaphorical meme dataset MET-Meme (Xu et al., 2022) has played a crucial role in advancing this field by catalyzing methodological innovations. The methods (He et al., 2024a, b; Wang et al., 2024) mainly focus on fine-grained feature alignment and fusion, resulting in suboptimal performance, as they often fail to fully leverage the characteristics of multimodal metaphors, overlooking the implied meanings and unconventional expressions that frequently appear in figurative tasks. On the other hand, methods (Xu et al., 2024b; Zhang et al., 2024; Xu et al., 2024a) that focus on generative knowledge expansion help bridge the gap between literal and figurative tasks by using the generative information from large language models (LLMs) or text-to-image models. Although using LLMs or text-to-image models has demonstrated significant advantages in multimodal metaphor identification tasks, they still face issues with high computational overhead and large GPU memory usage during training, even when these methods employ some general Parameter-Efficient Fine-Tuning (PEFT) techniques like LoRA (Hu et al., 2021). Recently, the novel methods (Zhou et al., 2023; Yu et al., 2024) that only fine-tune the LayerNorm layers of language models for cross-modal feature extraction and fusion, have demonstrated outstanding efficiency: achieving good performance by fine-tuning less than 4% of the total parameters. However, LayerNorm tuning of pretrained language models for feature extraction and fusion has remained unexplored in multimodal metaphor identification tasks due to their suboptimal performance when handling non-sequential data, such as images.
@@ 2.1. Multimodal Metaphor Understanding
The study of metaphors has expanded into the multimodal domain, exploring how this cognitive phenomenon manifests in various modalities, particularly in vision. Several datasets have been created to facilitate research in this area. MultiMET (Zhang et al., 2021) is a multimodal metaphor dataset containing text-image pairs with annotations for metaphor occurrence, domain relations, and sentiments. MetaCLUE (Akula et al., 2023) is another dataset including four understanding tasks for multimodal metaphor. The MET-Meme (Xu et al., 2022) dataset provides a valuable resource for studying multimodal metaphors in the context of memes. MultiCMET (Zhang et al., 2023) considers metaphors from different cultures, extending the work of MultiMET to Chinese. IRFL (Yosef et al., 2023) is a relevant dataset including metaphor and idiom data, which focuses on the multiple choices task. MemeCap (Hwang and Shwartz, 2023) includes annotations such as image captions, titles, and metaphorical captions to explore the automatic interpretation and understanding of multimodal metaphors. (R et al., 2024) introduces a task of video metaphor description, addressing the gap in Vision-Language (VL) models’ ability to understand metaphors in video. Various models have been proposed for detecting and understanding multimodal metaphors. Early work focused on visual atypicality, which can be seen as a special kind of multimodal metaphor, such as (Guo et al., 2021) models contextual compatibility to detect persuasive atypicality. The subsequent works primarily use MET-Meme as the benchmark, with early studies (He et al., 2024a, b; Wang et al., 2024; Zheng et al., 2025) focusing on fine-grained multimodal feature fusion or alignment. Recently, there is a growing trend to leverage pre-trained generative models to expand the knowledge for metaphor understanding. For example, C4MMD (Xu et al., 2024b) is a compact framework that uses a Chain-of-Thought (CoT) method to extract and integrate knowledge from Multimodal Large Language Models (MLLMs) into smaller models for multimodal metaphor detection. CAMEL (Zhang et al., 2024) leverages the captions generated by multimodal language models as a bridge to capture the implicitly established metaphor alignment. And MMMC (Xu et al., 2024a) utilizes a text-conditioned generative adversarial network to generate visual characteristics based on the linguistic attributes of metaphorical concepts, thereby capturing more comprehensive visual associations. Lately, ImaRA (Tian et al., 2025) leverages Retrieval-Augmented Generation to enhance LLMs’ capability to identify and interpret multimodal metaphors.
@@ 3.2. Concept Drift and Feature Fusion
[…]
@@ 4. Experiments
$\text{MET}\_\text{add}$ (Xu et al., 2022) 24.58 67.28 40.35 81.41 24.65 68.39 40.32 81.33
$\text{MET}\_\text{cat}$ (Xu et al., 2022) 28.04 66.70 38.87 82.54 27.68 67.25 38.56 82.39
@@
$\text{M}^{\text{3}}\text{F}\_\text{add}$ (Wang et al., 2024) 31.89 72.65 43.11 84.91 30.47 76.17 44.40 83.98
$\text{M}^{\text{3}}\text{F}\_\text{cat}$ (Wang et al., 2024) 31.85 71.77 44.33 84.56 29.82 74.09 44.10 83.20
MGMCF (Zheng et al., 2025) 35.98 77.92 47.72 88.04 34.36 78.11 47.92 87.51
@@
[…]
MultiCMET (Zhang et al., 2023) 78.79 85.66
VIEMF (He et al., 2024a) 83.92 84.87
SC-Net (He et al., 2024b) 86.85 87.50
C4MMD (Xu et al., 2024b) 79.05 88.04
ImaRA-7B (Tian et al., 2025) 83.45 91.04
CDGLT (Ours) 88.84 91.38
Table 2. Comparison of Macro F1-score on the Metaphor Identification (MI). The best scores are marked in bold.
@@ 4.2. Evaluation Metrics
(17) $\text{W-F1}=\frac{(2*Precision_{weighted}*Recall_{weighted})}{(Precision_{weighted}+Recall_{weighted})}$
@@ 4.4. Main Results
In this section, we compare our method with several existing multimodal metaphor detection methods, including MET-Meme (Xu et al., 2022), CAMEL (Zhang et al., 2024), $\text{M}^{\text{3}}\text{F}$ (Wang et al., 2024), MGMCF (Zheng et al., 2025), MultiCMET (Zhang et al., 2023), VIEMF (He et al., 2024a), SC-Net (He et al., 2024b), C4MMD (Xu et al., 2024b), and ImaRA (Tian et al., 2025). Among these methods, MultiCMET, VIEMF, SC-Net, C4MMD and ImaRA focus exclusively on the metaphor identification (MI) task and report macro-averaged F1 scores. The remaining methods are evaluated on four tasks in the MET-Meme dataset: Sentiment Analysis (SA), Offensiveness Detection (OD), Intention Detection (ID) and Metaphor Identification (MI). At the same time, MET-Meme, $\text{M}^{\text{3}}\text{F}$ and CAMEL have all proposed two model variants, and we have also included these variants in our comparison. As shown in Table 1 and Table 2, we report the performance results of the methods to be compared and our model settings. One of our settings is without SLERP called $\text{CDGLT}_{Vanilla}$ and the other with SLERP just called CDGLT. Although our approach is primarily focused on the multimodal metaphor identification task, we found that our architecture, with specifically training on other tasks, also achieved performance superior to or on par with the state-of-the-art. It is noted that CDGLT achieves the highest accuracy and Weighted F1 score on MI and SA tasks. On the other hand, $\text{CDGLT}_{Vanilla}$ demonstrates the complementary situations, achieving the better performance on the other two tasks: ID and OD. This suggests that incorporating Concept Drift embedding enhances the model’s ability to capture non-literal understanding tasks, such as MI. In contrast, tasks like ID and OD appear to benefit more from the vanilla model without SLERP, suggesting that these tasks demand more straightforward information and less divergence. This implies that the benefits of SLERP are not universally applicable and are instead highly dependent on the specific task. Our ablation analysis in section 4.5.1 will further clarify this situation. In addition, it can be observed that even with the use of Concept Drift, the CDGLT model achieves excellent performance in nearly all tasks, outperforming previous methods except in the accuracy of the OD task. Our framework does not utilize the generative information of LLM, surpassing methods like C4MMD, CAMEL and ImaRA, which use pre-trained language models to generate explanations of metaphors or implicit information as supplementary knowledge.
@@ 4.5. Further Analysis
In this section, we conducted a series of ablation experiments on the composition of $E^{Meme}$ and observed some interesting phenomena. As shown in Table 3, the results compare the input of the FFN adapter, $E^{Meme}$, with different settings. First, comparing $E^{I}$ only” versus $E^{T}$ only”, the results of the two experiments show only minor differences in the OD and ID tasks, while there are significant differences in the SA and MI tasks. Combined with other experiments, it is clear that the image modality is indispensable for improving performance across almost all tasks. This aligns with our intuition, as memes themselves use images as carriers of all information, including text. However, interestingly, for ID and OD tasks, just the OCR text from memes, although not the highest, still performs quite well. At the same time, this also suggests that, compared to the other two tasks, there may be more compatible commonalities between SA and MI, and the same applies to OD and ID.
The two experiments, $E^{S}$ only” and “concat($E^{I}$, $E^{T}$)”, reveal that on the SA and MI tasks, the performance of “concat($E^{I}$, $E^{T}$)” closely aligns with that of $E^{I}$ only”. Meanwhile, the performance of $E^{S}$ only” just falls between that of $E^{I}$ only” and $E^{T}$ only”. This may suggest that the concatenated feature vector aggregates information from both embeddings. However, the $E^{S}$ obtained through the SLERP operation seems to represent an intermediate state transitioning from $E^{I}$ to drift towards $E^{T}$.
@@
However, for the ID and OD tasks, there is a phenomenon where performing concatenation between any two embeddings of $E^{I}$, $E^{T}$, and $E^{S}$, results in improved performance on the ID or OD tasks compared to using $E^{I}$ or $E^{T}$ alone. Specifically, “concat($E^{I}$, $E^{S}$)” improves performance on the OD task compared to using $E^{I}$ only”, while “concat($E^{T}$, $E^{S}$)” enhances performance on the ID task compared to $E^{T}$ only”. Finally, “concat($E^{I}$, $E^{T}$)” leads to improvements in both the ID and OD tasks. Similar to “concat($E^{I}$, $E^{T}$)”, using $E^{S}$ only” also leads to improvements in both the ID and OD tasks compared to using $E^{I}$ only” and $E^{T}$ only”. Additionally, from the results of the group “concat($E^{I}$, $E^{S}$, $E^{T}$)” and “concat($E^{I}$, $E^{T}$)”, it can be observed that while adding $E^{S}$ improves the performance of SA and MI tasks, it also leads to a decline in the performance of ID and OD tasks. Furthermore, in the ID task, “concat($E^{I}$, $E^{S}$, $E^{T}$)” performs worse than “concat($E^{T}$, $E^{S}$)”, and in the OD task, it performs worse than “concat($E^{I}$, $E^{S}$)”. Therefore, concatenating any two embeddings with a third embedding will lead to a decrease in performance on both the OD and ID tasks. This could mean that the diverging information might be redundant for both OD and ID tasks, or even become noise that interferes with performance.
Comparison of the two experiments “concat($E^{I}$, $E^{S}$)” and “concat($E^{I}$, $E^{T}$)”, it shows that although replacing $E^{T}$ with $E^{S}$ leads to poorer performance on other tasks, it results in an improvement on the MI task. Additionally, by comparing “concat($E^{I}$, $E^{S}$)” with $E^{I}$ only” or “concat($E^{I}$, $E^{S}$, $E^{T}$)” with ’concat($E^{I}$, $E^{T}$)’, it is found that adding $E^{S}$ also contributes positively to the performance on the MI task. At the same time, this may suggest that divergent guidance, such as Concept Drift, indeed contributes to figurative tasks.
@@
Table 4. Prompt Ablations on Metaphor Identification (MI) task. “Visual Tokens” represent all the CLIP vision token features. $E^{I}$ represents the “CLS” token of CLIP vision encoder feature. Both are processed by a FFN adapter before being fed into GPT-2. F is the feature from concatenating $E^{T}$, $E^{S}$, and $E^{I}$ into an embedding three times the hidden dimension, then processed by the FFN adapter. The “length” column represents the number of embeddings of the sequence. Instruction A: “Providing you with the memes, now identify the metaphorical ones:”; Instruction B: “Metaphor Detection (literal — metaphor):”; Instruction C: “Metaphor Detection:”. Trainable and Frozen Vectors refer to prompts constructed using trainable and frozen Xavier initialization vectors, respectively.
@@
In this section, we conduct a series of ablation studies on the prompt construction method we proposed for LN Tuning. Experimental results on the MI task is shown in Table 4 demonstrating that, compared to other strategies, our method has significant advantages on the multimodal metaphor identification task based on meme data. The method of LN Tuning with a pretrained language model is commonly used for feature extraction or feature fusion of sequence information. Therefore, the most intuitive and common approach is to directly use all image patches (vision tokens) as the input to the language model. However, we found that by first fusing the multimodal global features (“CLS” tokens) using an FFN adapter and then directly using the fused features $F\in\mathbb{R}^{N\times 1\times d_{g}}$ as input, we achieved better performance, even though the input sequence length was 1 at that point. Since the attention mechanism exhibits stronger capabilities when processing sequences, we additionally construct an embedding sequence as a prompt based on $F$. We tried three strategies: the first one uses the word embedding layer of GPT2 to transform fixed instruction words into an embedding sequence; the second one uses Xavier initialization to create some learnable vectors; the third one freezes the Xavier-initialized vectors. In each method, $F$ is placed at the end of the sequence. For the first strategy, we tried three instructions: “Providing you with the memes, now identify the metaphorical ones:”, “Metaphor Detection (literal — metaphor):” and “Metaphor Detection:”. After tokenization, they consist of 14 tokens, 10 tokens and 5 tokens, respectively. For both “Trainable vectors” and “Frozen vectors”, we also used sequence lengths of 14, 10 and 5, respectively. The experimental results show that “frozen vectors” perform significantly better than “trainable vectors”. Comparing the experimental results of the Frozen Xavier initialization vectors method under different length settings may indicate that using prompts that are too long or too short could potentially lead to a decline in performance. However, we intuitively believe that when using word instructions as prompts, the performance will be significantly affected by the semantic content, and this impact is difficult to capture. At the same time, we noticed that in our experiments, the method of frozen vectors and word instructions performed similarly. To ensure the generality of the method, our framework ultimately chose to construct prompts by adopting frozen vectors.